OpenStudy (anonymous):

if the mean is subtracted from every number in a data set, and the differences are used to form a new data set, the data set must have a mean

OpenStudy (anonymous):

Equal to original mean Equal to zero Smaller than original mean Larger than original mean CAnnot be determined

OpenStudy (ganpat):

1 2 3 4 5 6 7 mean = 28/7 = 4 now subtracting 4 from all the numbers.. -3 -2 -1 0 1 2 3 4 so the new mean = ????